# Thermal Comfort in Sight: Thermal Affordance and its Visual Assessment for Sustainable Streetscape Design

### v1

@@ 3.1 Proposed definition of thermal affordance
As Gibson’s theory of affordance suggests (Gibson, 1977), environments contain inherent values and information that shape human perceptions and behaviours. Thermal comfort, as part of human perceptions, reflects both subjective satisfaction and objective factors such as air temperature, air humidity, and wind speed (Fanger et al., 1970; Nikolopoulou and Steemers, 2003; Van Hoof et al., 2010; Nicol and Roaf, 2017; de Feijter and van Vliet, 2021). We introduce the concept of ’thermal affordance’ to describe the inherent capability of an environment to impact thermal comfort. This concept integrates various environmental factors, indicating the possible thermal comfort experienced. Studies have shown the connection between thermal comfort and integrated environmental factors such as walkability (Labdaoui et al., 2021) and streetscape design (Yilmaz et al., 2021b), supporting the validity of thermal affordance.
Currently, the literature on thermal comfort presents various interconnected concepts, each providing a different lens to understand how individuals experience thermal environments (Lai et al., 2020). For instance, thermal comfort is defined by the ASHRAE Standard 55 as a ’condition of mind that expresses satisfaction with the thermal environment’, emphasising a subjective evaluation (ANSI/ASHRAE Standard 55, 2023). Meanwhile, thermal neutrality refers to the conditions where the majority of individuals feel neither warm nor cold (Shooshtarian and Rajagopalan, 2017; ANSI/ASHRAE Standard 55, 2023). Concepts such as the thermoneutral zone and thermal sensation further highlight the physiological and psychological dimensions of thermal comfort (Bligh and Johnson, 1973), with the latter focussing on how people consciously perceive temperature (Parsons, 2007; ANSI/ASHRAE Standard 55, 2023).
These established concepts offer significant insights, but they remain largely subjective, relying heavily on individuals’ preferences or perceptions. In contrast, our proposed notion of thermal affordance aims to transcend individual subjectivity by focussing on the inherent characteristics of the environment itself and its potential to influence thermal comfort. While more concepts like thermal expectation or thermal satisfaction consider personal inclinations and thermal experiences (De Dear and Schiller Brager, 2001; Nikolopoulou and Steemers, 2003; Dzyuban et al., 2022), thermal affordance unifies objective environmental features, offering a more scalable and integrative framework for OTC improvement in urban planning and design.
@@
  • Unity indicates that thermal affordance aims to encompass all fixed variables that influence thermal comfort in the environment into a unified set for description, based on the integration and unity properties of affordance proposed by Maier et al. (2009).
  • Objectivity highlights the objective environmental information encapsulated in thermal affordance, unaffected by individuals’ subjective perspectives, aligning with the idea of affordance as an integration of objectivity in the environment (Maier and Fadel, 2001).
  • Heuristic value of thermal affordance suggests that the information it encompasses can inspire understanding and analysis of thermal comfort, and the various dimensions of information within thermal affordance can mutually infer and enlighten one another. For instance, Ponce et al. (2018) suggests the possibility that visual data serve as heuristics for thermostats.
  • Spatial dependency highlights variations in thermal comfort caused by different spatial environments. Differences in spatial or environmental properties have been shown to directly lead to a difference in affordance (Fusco, 2016). Thermal affordance does not stress absolute levels of thermal comfort, but rather emphasises spatial differences between environments in their capacity to promote thermal comfort.
  • Interpretability means that thermal affordance has a clear inference relationship with environmental attributes, as environmental affordance is composed of multiple influencing parameters (Aradi et al., 2016). Comprehending this inference relationship can aid urban planners in enhancing thermal comfort through urban design.
  • Expandability means that thermal affordance can continuously incorporate additional environmental variables to update information on an environment’s capacity to influence thermal comfort. More dimensions of information, such as lifestyle and dwelling mode (Fusco, 2016), have been shown to allow for a more comprehensive evaluation and analysis of environmental affordance.
@@ 3.2 Visual assessment of thermal affordance (VATA) framework
This study identifies four key challenges in urban-scale OTC evaluation (Figure 2): (1) Measurement of thermal comfort requires substantial resources – labour, budget, and time (Kumar and Sharma, 2020). (2) Current methods, relying on physical instruments and surveys, can only cover limited urban areas (Balogun and Daramola, 2019). (3) Geographic precision is constrained for city-wide OTC assessments (Mutani et al., 2022). (4) Urban OTC measurements are affected by many uncontrollable environmental variables, unlike controlled indoor experiments (Chen and Ng, 2012). To address these challenges, we developed the VATA (visual assessment of thermal affordance) framework to enhance urban-scale OTC assessments and improve understanding of urban thermal environments.
@@
This research continues to explore the intrinsic connection between OTC in urban environments and visual data from SVI, shaped by affordance information and personal experiences. First, SVI visual data reflect objective characteristics of streetscapes, indicating streetscapes’ potential capability to promote thermal comfort and microclimate conditions (Qaid et al., 2018; Yilmaz et al., 2021a), which forms thermal affordance. Second, this visual information is linked to individuals’ thermal perception, influenced by memory and sensory experiences (Ko et al., 2020; Yang et al., 2021; Nitidara et al., 2022), which affects their thermal affordance assessment. These insights form the foundation of our VATA analysis, highlighting the role of visual characteristics and perceptual experiences in OTC assessment. Using machine learning, we combine SVI image features (IF) and visual-perceptual indicator (VPI) survey data to predict the VATA metric, allowing us to visually assess the urban environments’ capability to promote thermal comfort. Our model shows promise for streetscape assessment and VATA prediction, as illustrated in Figure 2.
@@
Objective IF from SVI reflect the urban environment from an individual’s perspective in a three-dimensional, objective manner (Biljecki and Ito, 2021; Hou and Biljecki, 2022), capturing elements such as greening rate (Li et al., 2015), building information (Kang et al., 2018), urban function (Zhang et al., 2023), and urban form (Biljecki et al., 2023). These features, closely related to OTC, serve as affordance information since OTC is influenced by surrounding urban elements (Yang et al., 2011; Deng et al., 2023), enabling visual-environmental inference in urban studies (Kang et al., 2020; Kruse et al., 2021; Guan et al., 2022). The VPI information obtained through the SVI visual assessment surveys further accounts for urban perceptions of individuals (Ma et al., 2021; Ito et al., 2024a), impacting socioeconomic dimensions such as bikeability (Ito and Biljecki, 2021), running ability (Dong et al., 2023) and property prices (Yang et al., 2023). VPI and OTC, as psychological perceptions, subtly influence each other (Klemm et al., 2015; Shooshtarian and Ridley, 2017; Lam et al., 2020; Liu et al., 2024). Figure 1 illustrates the relationships between IF, VPI, VATA and OTC, demonstrating that the use of SVI for the evaluation of VATA and OTC is grounded in both environmental and perceptual principles.
@@
From another perspective, the intricate relationship between VATA and OTC can be deeply rooted in individuals’ long-term memory and experiential knowledge, as people’s sensory experience in urban environments is a long-term process. This kind of environmental experience goes beyond short-term observations and involves complex reasoning and memory processes in the brain (Knauff, 2009; Öğmen and Herzog, 2016; Wang et al., 2021). For example, the association of greenery with shading and cooling stems from prolonged human experiences (Bar, 2009; Eichenbaum and Fortin, 2009). This knowledge enables individuals to infer OTC from visual cues, similar to how greenery is linked to cooler temperatures and better thermal comfort(Friston, 2003; Berkman and Falk, 2013; Ma et al., 2024). Thus, subjective streetscape visual assessments in our VATA framework are also grounded in the brain’s ability to interpret environmental cues based on past experiences.
@@ 4.1 Research framework and study area
Figure 3 presents the research framework for this study, built on the VATA framework. We conducted an online SVI visual assessment survey, evaluating VATA and 19 other VPIs based on 500 SVIs. Five classes of IFs from SVI data, along with survey-based VATA and VPI data, were used to develop datasets for statistical VATA prediction and inference models. A multi-task neural network learning model (MTNNL) was constructed with two stages: predicting VPIs from IFs, and then predicting VATA from VPIs, using weighted loss values for iterative training. This model was applied to predict VATA for 92,233 SVIs in Singapore, and validated against real-world OTC data. A two-stage elastic net regression model (ENRM) was also used to interpret IF-VPI-VATA relationships, offering insights for streetscape design. SVIs are sourced from Google Street View (Anguelov et al., 2010).
Singapore serves as the study area due to its rich urban landscape design, diverse urban forms, consistent outdoor temperatures (25-31 °C year-round) and perennial need to ensure OTC as a tropical city (Yang et al., 2013). Our study includes an online SVI visual assessment survey, model prediction and inference, and validation using OTC field data, aiming to propose a more efficient method for urban-scale OTC assessment.
@@ 4.2 Online SVI assessment survey setting
Along with VATA, the 19 VPIs are grouped into four categories: microclimate inference, built environment evaluation, streetscape design quality, and evoked emotion (Table 1). Participants were presented with system-randomised pairs of SVIs from the online survey for each indicator, which they then compared. The TrueSkill algorithm (Herbrich et al., 2006) was subsequently employed to analyse responses, generating rankings and scores for the 500 SVIs on VATA and each VPI. These scores were then normalised to a 0-5 scale for machine learning regression (Figure 4). The ranking algorithm consistently updates the ranking score of winners and losers from pairwise comparisons, utilising the survey results comprising 3168 pairs collected from all the 176 participants for scoring calculation. Since one assumption of the TrueSkill algorithm is that the performance of all ranking elements (500 images in this case) follows a normal distribution, we can ultimately visualise the scoring results as a normal distribution within a range of 0 to 5 after normalisation, with the ideal population mean of 2.5. The TrueSkill process was repeated 20 times to obtain scores for VATA and each VPI. Details of the TrueSkill algorithm and survey results are introduced in A and B.
To enhance model training based on SVIs and corresponding indicator scores from the online SVI visual assessment survey, we apply six data augmentations—resizing, normalising, random clipping, rotating, color jittering, and cropping—to the 500 images. These transformations enhance the model’s generalisation and robustness (Figure 4). For VATA and VPI scores, we set the standard deviation (std) of each scoring distribution to 1, balancing discriminative power and an even score distribution. As illustrated in Figure 4, a std of 0.5 centralises the scores, limiting range, while a std of 1.5 flattens the distribution, reducing distinction. Setting std to 1 strikes an optimal balance between clarity and evenness.
@@
    k-means clustering of all SVIs based on 19 streetscape elements from semantic segmentation, following the streetscape classification method outlined by Liang et al. (2023).
@@ 4.3 IF extracted from SVI data for model training and prediction
As illustrated in Table 2, five categories of IF can be extracted from a single SVI $i$ using various algorithms and datasets: semantic segmentation identifies streetscape elements (e.g., buildings, vegetation), convolutional data capture high-dimensional pixel representations, object recognition counts objects (e.g., people, cars), pixel feature analysis examines colour features, and scene recognition assesses the likelihood of displaying specific scene type (e.g., commercial, natural). Pretrained models like DeepLabV3+ (Chen et al., 2018), ResNet-50 (He et al., 2016), Faster R-CNN (Ren et al., 2015), and OpenCV (Bradski, 2000), along with datasets like Cityscape (Cordts et al., 2016), COCO (Lin et al., 2014), and Places365 (Zhou et al., 2017a), are used to extract these IFs.
@@ 4.4 Architecture of VATA prediction and inference model
VATA employs two main model architectures: the multi-task neural network learning (MTNNL) for predictive analysis and the elastic net regression model (ENRM) for inference analysis. MTNNL is used for predicting VATA outcomes based on SVI and human visual assessment results, leveraging its strong predictive power. ENRM, as a linear and interpretable model, addresses collinearity issues and is suited for VATA inference research.
@@ 4.5 VATA validation based on OTC dataset
To validate the effectiveness of the VATA framework in assessing human thermal comfort in urban street environments, we employed an on-the-ground OTC field survey dataset and data-driven statistical evaluations. The validation utilised OTC measurements from field investigations conducted on the campus of the National University of Singapore (NUS) (Liu et al., 2023).
The survey involved 15 NUS students and 43 investigation points along a continuous path (Figure 7). The participants walked the path, providing comfort ratings (1-10 scale) at each point. They wore Apple Watches with the Cozie app (Tartarini et al., 2023) to collect subjective feedback and heart rate data. Sound and solar intensity were measured using a sound metre and a smartphone, respectively. These ’thermal walk’ experiments (Ignatius et al., 2024) were carried out three times daily (10 am, 2 pm, 5 pm) over five consecutive days in July 2022. The resulting OTC dataset includes average comfort scores, heart rate values, sunlight intensity, noise levels, and altitude measurements.
In addition, Google Street View images corresponding to each of the 43 investigation points were collected to facilitate visual-based VATA predictions and validations. The validation methodology, described in Figure 7, comprises two parts: 1) comparing VATA with OTC to assess thermal affordance effectiveness by directly comparing SVI-predicted VATA at 43 data points with corresponding OTC values; and 2) using a multivariate linear modelling approach combining VATA and HSNA indicators (heart rate, solar intensity, noise level, and altitude) to evaluate VATA reliability in predicting OTC.

### v4

@@ 2.1 Proposed definition of thermal affordance
As Gibson’s theory of affordance suggests [78], environments contain inherent values and information that shape human perceptions and behaviours. Thermal comfort, as part of human perceptions, reflects both subjective satisfaction and objective factors such as air temperature, air humidity, and wind speed [24, 107, 108, 109, 110]. We introduce the concept of ’thermal affordance’ to describe the inherent capability of an environment to impact thermal comfort. This concept integrates various environmental factors, indicating the possible thermal comfort experienced. Studies have shown the connection between thermal comfort and integrated environmental factors such as walkability [97] and streetscape design [75], supporting the validity of thermal affordance.
Currently, the literature on thermal comfort presents various interconnected concepts, each providing a different lens to understand how individuals experience thermal environments [111]. For instance, thermal comfort is defined by the ASHRAE Standard 55 as a ’condition of mind that expresses satisfaction with the thermal environment’, emphasising a subjective evaluation [112]. Meanwhile, thermal neutrality refers to the conditions where the majority of individuals feel neither warm nor cold [113, 112]. Concepts such as the thermoneutral zone and thermal sensation further highlight the physiological and psychological dimensions of thermal comfort [114], with the latter focussing on how people consciously perceive temperature [115, 112].
These established concepts offer significant insights, but they remain largely subjective, relying heavily on individuals’ preferences or perceptions. In contrast, our proposed notion of thermal affordance aims to transcend individual subjectivity by focussing on the inherent characteristics of the environment itself and its potential to influence thermal comfort. While more concepts like thermal expectation or thermal satisfaction consider personal inclinations and thermal experiences [116, 107, 117], thermal affordance unifies objective environmental features, offering a more scalable and integrative framework for OTC improvement in urban planning and design.
@@
  • Unity indicates that thermal affordance aims to encompass all fixed variables that influence thermal comfort in the environment into a unified set for description, based on the integration and unity properties of affordance proposed by [118].
  • Objectivity highlights the objective environmental information encapsulated in thermal affordance, unaffected by individuals’ subjective perspectives, aligning with the idea of affordance as an integration of objectivity in the environment [119].
  • Heuristic value of thermal affordance suggests that the information it encompasses can inspire understanding and analysis of thermal comfort, and the various dimensions of information within thermal affordance can mutually infer and enlighten one another. For instance, [120] suggests the possibility that visual data serve as heuristics for thermostats.
  • Spatial dependency highlights variations in thermal comfort caused by different spatial environments. Differences in spatial or environmental properties have been shown to directly lead to a difference in affordance [80]. Thermal affordance does not stress absolute levels of thermal comfort, but rather emphasises spatial differences between environments in their capacity to promote thermal comfort.
  • Interpretability means that thermal affordance has a clear inference relationship with environmental attributes, as environmental affordance is composed of multiple influencing parameters [121]. Comprehending this inference relationship can aid urban planners in enhancing thermal comfort through urban design.
  • Expandability means that thermal affordance can continuously incorporate additional environmental variables to update information on an environment’s capacity to influence thermal comfort. More dimensions of information, such as lifestyle and dwelling mode [80], have been shown to allow for a more comprehensive evaluation and analysis of environmental affordance.
@@ 2.2 Visual assessment of thermal affordance (VATA) framework
This study identifies four key challenges in urban-scale OTC evaluation (Figure 2): (1) Measurement of thermal comfort requires substantial resources – labour, budget, and time [122]. (2) Current methods, relying on physical instruments and surveys, can only cover limited urban areas [123]. (3) Geographic precision is constrained for city-wide OTC assessments [124]. (4) Urban OTC measurements are affected by many uncontrollable environmental variables, unlike controlled indoor experiments [10]. To address these challenges, we developed the VATA (visual assessment of thermal affordance) framework to enhance urban-scale OTC assessments and improve understanding of urban thermal environments.
@@
This research continues to explore the intrinsic connection between OTC in urban environments and visual data from SVI, shaped by affordance information and personal experiences. First, SVI visual data reflect objective characteristics of streetscapes, indicating streetscapes’ potential capability to promote thermal comfort and microclimate conditions [125, 126], which forms thermal affordance. Second, this visual information is linked to individuals’ thermal perception, influenced by memory and sensory experiences [127, 128, 129], which affects their thermal affordance assessment. These insights form the foundation of our VATA analysis, highlighting the role of visual characteristics and perceptual experiences in OTC assessment. Using machine learning, we combine SVI image features (IF) and visual-perceptual indicator (VPI) survey data to predict the VATA metric, allowing us to visually assess the urban environments’ capability to promote thermal comfort. Our model shows promise for streetscape assessment and VATA prediction, as illustrated in Figure 2.
@@
Objective IF from SVI reflect the urban environment from an individual’s perspective in a three-dimensional, objective manner [35, 130], capturing elements such as greening rate [131], tree canopies[132], building information [65], urban function [133], and urban form [134]. These features, closely related to OTC, serve as affordance information since OTC is influenced by surrounding urban elements [135, 136], enabling visual-environmental inference in urban studies [42, 82, 137]. The VPI information obtained through the SVI visual assessment surveys further accounts for urban perceptions of individuals [138, 56], impacting socioeconomic dimensions such as bikeability [90], running ability [139] and property prices [70]. VPI and OTC, as psychological perceptions, subtly influence each other [74, 140, 141, 142]. Figure 1 illustrates the relationships between IF, VPI, VATA and OTC, demonstrating that the use of SVI for the evaluation of VATA and OTC is grounded in both environmental and perceptual principles. These relationships are not directly derived or scored from the survey data, as the problem remains a “black box” with only separate scores available for IF, VPI, VATA, and OTC. Accordingly, machine learning techniques—including both predictive and inferential modelling—are used to uncover and quantify these underlying relationships.
@@
From another perspective, the intricate relationship between VATA and OTC can be deeply rooted in individuals’ long-term memory and experiential knowledge, as people’s sensory experience in urban environments is a long-term process. This kind of environmental experience goes beyond short-term observations and involves complex reasoning and memory processes in the brain [143, 144, 145]. For example, the association of greenery with shading and cooling stems from prolonged human experiences [146, 147]. This knowledge enables individuals to infer OTC from visual cues, similar to how greenery is linked to cooler temperatures and better thermal comfort[148, 149, 150]. Thus, subjective streetscape visual assessments in our VATA framework are also grounded in the brain’s ability to interpret environmental cues based on past experiences.
While VATA provides an efficient means of quantifying thermal affordance using SVI data, it is important to acknowledge that integrating meteorological information, detailed geospatial data, remote sensing imagery, and socio-cultural contexts can yield more comprehensive insights into how multiple factors collectively influence urban thermal comfort [27, 151, 152]. Incorporating additional environmental variables could further enhance the VATA framework by employing advanced data fusion and machine learning techniques, ultimately improving predictive accuracy and informing more effective urban design strategies.
@@ 3.1 Research framework and study area
Figure 3 presents the research framework for this study, built on the VATA framework. We conducted an online SVI visual assessment survey, evaluating VATA and 19 other VPIs based on 500 SVIs. Five classes of IFs from SVI data, along with survey-based VATA and VPI data, were used to develop datasets for statistical VATA prediction and inference models. Using machine learning models, we can predict and infer relationships between IFs, VPIs, and VATA based on their respective scoring data, all drawn from the same set of 500 SVIs used in the visual assessment survey. A multi-task neural network learning model (MTNNL) was constructed with two stages: predicting VPIs from IFs, and then predicting VATA from VPIs, using weighted loss values for iterative training. This model was applied to predict VATA for 92,233 SVIs in Singapore, and validated against real-world OTC data. A two-stage elastic net regression model (ENRM) was also used to interpret IF-VPI-VATA relationships, offering insights for streetscape design. SVIs are sourced from Google Street View [153].
Singapore serves as the study area due to its rich urban landscape design, diverse urban forms, consistent outdoor temperatures (25-31 °C year-round) and perennial need to ensure OTC as a tropical city [154]. Our study includes an online SVI visual assessment survey, model prediction and inference, and validation using OTC field data, aiming to propose a more efficient method for urban-scale OTC assessment.
@@ 3.2 Online SVI assessment survey setting
Along with VATA, the 19 VPIs are grouped into four categories: microclimate inference, built environment evaluation, streetscape design quality, and evoked emotion (Table 1). These VPIs were chosen to capture participants’ holistic visual-perceptual impressions of streetscapes, rather than having them rate individual streetscape elements. Participants were presented with system-randomised pairs of SVIs from the online survey for each indicator, which they compared directly. This approach encouraged a more natural expression of their perceptions, free from the pressure of explicit grading tasks. The TrueSkill algorithm [155] was subsequently employed to analyse responses, generating rankings and scores for the 500 SVIs on VATA and each VPI. These scores were then normalised to a 0-5 scale for machine learning regression (Figure 4). The ranking algorithm consistently updates the ranking score of winners and losers from pairwise comparisons, utilising the survey results comprising 3168 pairs collected from all the 176 participants for scoring calculation. Since one assumption of the TrueSkill algorithm is that the performance of all ranking elements (500 images in this case) follows a normal distribution, we can ultimately visualise the scoring results as a normal distribution within a range of 0 to 5 after normalisation, with the ideal population mean of 2.5. The TrueSkill process was repeated 20 times to obtain scores for VATA and each VPI. Details of the TrueSkill algorithm and survey results are introduced in A and B. These TrueSkill-derived scores enable us to employ predictive and inferential modelling to construct the IF–VPI–VATA relationship.
To enhance model training based on SVIs and corresponding indicator scores from the online SVI visual assessment survey, we apply six data augmentations—resizing, normalising, random clipping, rotating, colour jittering, and cropping—to the 500 images. These transformations enhance the model’s generalisation and robustness (Figure 4). For VATA and VPI scores, we set the standard deviation (std) of each scoring distribution to 1, balancing discriminative power and an even score distribution. As illustrated in Figure 4, a std of 0.5 centralises the scores, limiting range, while a std of 1.5 flattens the distribution, reducing distinction. Setting std to 1 strikes an optimal balance between clarity and evenness.
@@
    $k$-means clustering of all SVIs based on 19 streetscape elements from semantic segmentation, following the streetscape classification method outlined by [37].
@@ 3.3 IF extracted from SVI data for model training and prediction
As illustrated in Table 2, five categories of IF can be extracted from a single SVI $i$ using various algorithms and datasets: semantic segmentation identifies streetscape elements (e.g., buildings, vegetation), convolutional data capture high-dimensional pixel representations, object recognition counts objects (e.g., people, cars), pixel feature analysis examines colour features, and scene recognition assesses the likelihood of displaying specific scene type (e.g., commercial, natural). Pre-trained models like DeepLabV3+ [156], ResNet-50 [157], Faster R-CNN [158], and OpenCV [159], along with datasets like Cityscape [59], COCO [160], and Places365 [161], are used to extract these IFs.
@@ 3.4 Architecture of VATA prediction and inference model
VATA employs two main model architectures: the multi-task neural network learning (MTNNL) for predictive analysis and the elastic net regression model (ENRM) for inference analysis. MTNNL is used for predicting VATA outcomes based on SVI and human visual assessment results, leveraging its strong predictive power. ENRM, as a linear and interpretable model, addresses collinearity issues and is suited for VATA inference research. With separate scores obtained for IFs, VPIs, and VATA for each image in the sample, we can use MTNNL to predict VPIs and VATA scores from IF data, and employ ENRM to quantitatively determine how the visual environmental features encapsulated in IFs influence both VPIs and VATA. In this process, VPIs act as an intermediate layer that not only improves VATA prediction accuracy but also deepens our understanding of the IF–VPI–VATA relationships.
@@ 3.5 VATA validation based on OTC dataset
To validate the effectiveness of the VATA framework in assessing human thermal comfort in urban street environments, we employed an on-the-ground OTC field survey dataset and data-driven statistical evaluations. The validation utilised OTC measurements from field investigations conducted on the campus of the National University of Singapore (NUS) [162]. These field investigations were carried out independently from both the SVI-based visual assessments and the subsequent VATA modelling efforts, ensuring that there was no pre-established alignment or dependency between the datasets. This relative independence enhances the robustness of the validation process by minimizing the risk of bias and reinforcing the credibility of the comparative analysis.
The survey involved 15 NUS students and 43 investigation points along a continuous path (Figure 7). The participants walked the path, providing comfort ratings (1-10 scale) at each point. They wore Apple Watches with the Cozie app [163] to collect subjective feedback and heart rate data. Sound and solar intensity were measured using a sound metre and a smartphone, respectively. These ’thermal walk’ experiments [164] were carried out three times daily (10 am, 2 pm, 5 pm) over five consecutive days in July 2022. The resulting OTC dataset includes average comfort scores, heart rate values, sunlight intensity, noise levels, and altitude measurements. We specifically focused on these parameters—readily captured via portable devices such as smartphones and wearable sensors—due to their potential for rapid variability across short distances. This enables fine-grained assessment of thermal comfort differences along the walking route, without the added complexity of factors like humidity or air temperature that tend to remain relatively uniform over these small spatial scales. While including additional factors such as humidity or air temperature measurements could enhance the comprehensiveness of comfort modelling, doing so is not the central focus of this study. Additionally, although Singapore provides publicly available air temperature data, the uneven spatial distribution and limited number of weather stations (where sensors are located) create significant uncertainties when interpolating localized conditions. Furthermore, the lack of measurements at very short intervals further reduces the granularity of these data for our fine-scale analysis.
In addition, Google Street View images corresponding to each of the 43 investigation points were collected to facilitate visual-based VATA predictions and validations. The Google Street View imagery used was captured in the same year as the field investigation (2022), thus ensuring temporal synchronization. This alignment in timing allowed for a more reliable comparison and modelling between the VATA prediction scores and the on-site OTC measurements. The validation methodology, described in Figure 7, comprises two parts: 1) comparing VATA with OTC to assess thermal affordance effectiveness by directly comparing SVI-predicted VATA at 43 data points with corresponding OTC values; and 2) using a multivariate linear modelling approach combining VATA and HSNA indicators (heart rate, solar intensity, noise level, and altitude) to evaluate VATA reliability in predicting OTC.